OpenStudy (anonymous):

find the value of h and k that makes the equation true x^2+4x= (x-h)^2 + k

OpenStudy (anonymous):

h=-2 k=-4

OpenStudy (anonymous):

or h=0 k=4x

OpenStudy (anonymous):

how'd you get the answer jahtoday?

OpenStudy (anonymous):

re-write your equation as: \[x ^{2}+4x=x ^{2}-2hx+h ^{2}+k\] then: \[4x=x* (-2h) +(k+h ^{2})\] in order left side = right one, we need what...?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

4=-2h , which gives: h=-2 and k+h^2=0, which will give you: k=-4

OpenStudy (anonymous):

oooh okay..